# FORCE: Transferable Visual Jailbreaking Attacks via Feature Over-Reliance CorrEction

### v2

@@ 4 Experiment
In this section, we evaluate the effectiveness of FORCE, including experimental setups (Section 4.1), performance evaluations (Section 4.2), and ablation studies (Section 4.3).
@@ 4.1 Experimental Setups
Source Models and Baselines. We use LLaVA-v1.5-7B (Liu et al., 2023a) as the source MLLM for both the baseline and our proposed method. For the baseline, we adopt standard PGD (Madry et al., 2018) to generate visual attacks, with a step size of $2/255$ and a perturbation budget of $32/255$. The optimisation target is set to Sure, here is. In this work, we consider two attack settings: zero-shot and multi-query. In the zero-shot setting, we only craft one visual attack that satisfies the optimisation objective on the source MLLM and then directly evaluate it on the target MLLMs. In the multi-query, we generate 100 distinct visual jailbreaking examples that meet the optimisation target on the source model and evaluate them individually on the target model.
@@
Target Models. We select a range of popular safety-aligned MLLMs as transfer-target, treating them as black-box models with inaccessible parameters. For adapter-based MLLMs, we use InstructBLIP-Vicuna-7B (Dai et al., 2023), Llava-v1.6-mistral-7b (Liu et al., 2023a), and Idefics3-8B-Llama3 (Laurençon et al., 2024). For early-fusion MLLMs, we evaluate Qwen2.5-VL-7B-Instruct (Bai et al., 2023b) and LLaMA-3.2-11B-Vision-Instruct (Meta, 2024). For commercial MLLMs, we consider Claude-Sonnet-4 (Anthropic, 2025), Gemini-2.5-Pro (Google, 2025), and GPT-5 (OpenAI, 2025).
@@
Datasets and Evaluation Metrics. We evaluate our approach on three benchmarks: MaliciousInstruct (Huang et al., 2024), AdvBench (Zou et al., 2023), and HADES (Li et al., 2024), containing 100, 520, and 750 malicious instructions, respectively. For textual inputs, we adopt plain malicious prompts without modification. For AdvBench and MaliciousInstruct, the visual input is initialised with either a blank image of RGB $(128,128,128)$ or a panda image (Qi et al., 2024). For HADES, we adopt the provided image–instruction pairs (step 5) as initialisation while removing keyword typography to ensure the model focuses on the image content. Regarding commercial models, we test the top 100 instructions from MaliciousInstruct and AdvBench, and the top 20 instructions in HADES spanning five categories. To avoid false positives, we evaluate the attack success rate (ASR) by combining substring matching with LLM-based judgment. Substring matching verifies whether the model refuses to answer the malicious instruction (Zou et al., 2023), while HarmBenchLLaMA-2-13B-cls (Mazeika et al., 2024) determines whether the response is actually harmful.
@@
Setup for FORCE. We set the number of reference samples $N=10$, the noise neighbourhood $\eta=8/255$, the regularisation strength $\lambda=1$, and the number of frequency bands $M=10$. All other settings remain consistent with the baseline PGD to ensure a fair comparison.
4.2 Performance Evaluation
To comprehensively evaluate our attack, we examine its cross-model transferability on two different MLLM architectures and API-based MLLMs. From Table 1, we can observe that visual jailbreaking attacks generated by standard PGD exhibit considerable transferability to adapter-based MLLMs, with an average ASR of about 50% and requiring 50 queries per successful attack. For this scenario, our proposed FORCE demonstrates superior performance across all evaluation settings, achieving an average ASR improvement of 13% while reducing the average query cost by over 10%.
Finally, we extend our method to jailbreak commercial MLLMs, which incorporate state-of-the-art alignment techniques and auxiliary safety filters. As shown in Table 1, FORCE can consistently enhance transferability across three mainstream commercial models, achieving an average improvement of 70%. Despite the baseline’s limited capability restricting absolute ASR increases, our method delivers substantial relative improvements and represents a firm step toward practical optimisation-based visual attacks. The real-world case analysis of FORCE attacks can be found in Appendix C.
@@ 4.3 Ablation Study
Zero-shot Transferability. We further report the most stringent zero-shot transferability, where only a single query is permitted to jailbreak the target MLLMs. From Table 2 (right), this restrictive scenario leads to a sharp decline in the PGD effectiveness, which can be attributed to its narrow feasible regions that are hard to precisely match with the vulnerabilities of target models. While this setting also poses challenges for FORCE, its ability to discover a flatter loss landscape increases the likelihood of exploiting target vulnerabilities with a single attempt and improves transferability.
@@
Optimisation Objectives. To validate the effectiveness of our proposed method in reducing model-specific reliance, we visualise the layers’ feasible regions and the influence of frequency bands. These visualisations follow the same approach described in Section 3.2 and Section 3.3. As presented in Figure 5, it is clear that our method encourages visual jailbreaking attacks to explore broader
@@
representations in the early layers, resulting in a smoother loss increase during feature interpolation compared to the baseline in Figure 3. This also drives the attack toward a flatter loss landscape in the input space, thereby improving its resilience to parameter shifts during transfer. We also examine the capability of our method in the spectral domain by analysing the influence of frequency components on attack performance. As depicted in Figure 6, our method reliably mitigates the abnormal reliance on semantically poor information, as evidenced by a more moderate loss change when masking high-frequency informations, and exhibits a natural trend similar to that of non-adversarial images. Both outcomes indicate that FORCE effectively mitigates model-specific reliance, promotes exploration of a flatter loss landscape, and enhances transferability.
Impact of Components. We investigate the individual and the synergistic impact of the two components of our algorithm, as presented in Table 3. Our results demonstrate that each component can effectively mitigate its targeted reliance. Specifically, transferability improves by 5.7% and query efficiency by 8.2% with the layer-feature regularisation, and by 15.1% and 18.8% with the spectral-rescaling component. The synergy of the two components enables a more thorough removal of improper reliance, resulting in an overall performance gain of 18.9%.
@@
$\checkmark$ - 56.00 (5.7%) 46.89 (8.2%)
- $\checkmark$ 61.00 (15.1%) 42.70 (18.8%)
$\checkmark$ $\checkmark$ 63.00 (18.9%) 41.18 (23.2%)
@@ 5 Conclusion
In this work, we investigated the limited transferability of optimisation-based visual jailbreaking attacks and attributed this issue to their reliance on model-specific features in early layers and high-frequency information. This reliance drives the attacks into high-sharpness regions, leaving them vulnerable to parameter shifts during transfer. To address this, we introduced a Feature Over-Reliance CorrEction (FORCE) method, which encourages attacks to explore broader regions in the layer space while rescaling frequency components according to their semantic relevance. By correcting both layer space and spectral domain dependencies, FORCE enables the discovery of flattened feasible regions that enhance cross-model transferability. Extensive experiments demonstrate that our approach provides an important step toward a practical visual red-teaming evaluation.
The Use of Large Language Models
During the preparation of this manuscript, the authors used LLMs to support the writing process, including polishing language, improving clarity, and enhancing overall readability. Since this work directly concerns the safety of MLLMs, we also employ them in our experiments. MLLMs are used to generate visual jailbreaking attacks, they are evaluated as victim models to test attack transferability, and they are employed as judges to determine whether the attacks are successful. The authors retained full intellectual control over this paper. LLMs were employed solely as tools to support the research process, and the final manuscript represents the authors’ original work and insights.
Ethics Statement
The authors acknowledge that the techniques presented in this paper could, in principle, be misused to maliciously manipulate MLLMs. However, our intention is to provide insights into the inherent weaknesses of MLLMs from a red-teaming perspective, with the goal of encouraging blue-team efforts to strengthen their safety mechanisms. We believe that transparent discussions of potential threats are fundamental to shaping human-aligned MLLMs, and we hope this work will promote safety, accountability, and ethical practices in their deployment.
Reproducibility Statement
We have made every effort to ensure the reproducibility of our work. Comprehensive details of the experimental settings, hyperparameters, and algorithms are provided in both the main paper and the appendix. All baselines, datasets and MLLMs used in this study are publicly available through GitHub, HuggingFace, or accessible APIs. All computations were conducted on AMD MI250X GPUs with 128 GB of memory, and can also be reproduced on GPUs with smaller memory capacity.
@@ Appendix A Feature Interpolation Between Visual Jailbreaking Attacks
We also interpolate features between two different visual jailbreaking attacks, as shown in Figure 7. Consistent with our observation in Figure 3, we find that feasible regions in later layers are flatter, whereas they become progressively narrower toward earlier layers. Moreover, our results show that in later layers, different jailbreaking examples occupy a shared continuous region, as interpolated attacks consistently succeed in manipulating the source MLLM. In earlier layers, the feasible regions of different attacks become disjoint, as the interpolated features cause them to lose effectiveness. Togetherly, these results reveal that visual attacks tend to rely on model-specific features in earlier layers, leading to small and disjoint feasible regions that fail to generalise across models.
@@
Appendix B Feature Over-Reliance CorrEction Algorithm
@@
3:  Generate spectral-rescaled perturbation via Algorithm Algorithm 3 $\delta\leftarrow\delta_{rescaled}$
@@
0: L-layer Network $f_{\mathbf{\theta}}$, input text $\mathbf{x}_{\text{txt}}$, input image $\mathbf{x}_{\text{img}}$, target output $\mathbf{y}$, jailbreaking perturbation $\delta$, number of reference samples $N$, noise neighbourhood $\eta$, regularisation strength $\lambda$.
@@
[…]
@@
4:  Extract layer feature $\mathbf{h}_{\eta_{n},l}=\!\left(f_{\theta,l}(\mathbf{x}_{\text{img}}+\delta+\eta_{n},\mathbf{x}_{\text{txt}})\right),\quad$ for $l=1,\ldots,L$
@@
7: Extract layer feature $\mathbf{h}_{jail,l}=\!\left(f_{\theta,l}(\mathbf{x}_{\text{img}}+\delta,\mathbf{x}_{\text{txt}})\right),\quad$ for $l=1,\ldots,L$
[…]
@@
0: L-layer Network $f_{\mathbf{\theta}}$, input text $\mathbf{x}_{\text{txt}}$, input image $\mathbf{x}_{\text{img}}$, target output $\mathbf{y}$, jailbreaking perturbation $\delta$, number of frequency bands $M$.
@@
[…]

### v3

@@ 4 Experiment
In this section, we evaluate the effectiveness of FORCE, including experimental setups (Section 4.1), performance evaluations (Section 4.2), ablation studies (Section 4.3), and generation costs (Section 4.4).
@@ 4.1 Experimental Setups
Source Models and Baselines. We use LLaVA-v1.5-7B [22] as the source MLLM for both the baseline and our proposed method. The results obtained from different source models are provided in Appendix E. For the baseline, we adopt standard PGD [26] to generate visual attacks, with a step size of $2/255$ and a perturbation budget of $32/255$. The optimisation target is set to ‘‘Sure, here is’’. In this work, we consider two attack settings: zero-shot and multi-query. In the zero-shot setting, we only craft one visual attack that satisfies the optimisation objective on the source MLLM and then directly evaluate it on the target MLLMs. In the multi-query, we generate 100 distinct visual jailbreaking examples that meet the optimisation target on the source model and evaluate them individually on the target model. We also compare our method with the textual jailbreaking attack GCG [64] in Appendix F, as well as with other visual attack baselines in Appendix G.
@@
Target Models. We select a range of popular safety-aligned MLLMs as transfer-target, treating them as black-box models with inaccessible parameters. For adapter-based MLLMs, we use InstructBLIP-Vicuna-7B [9], Llava-v1.6-mistral-7b [22], and Idefics3-8B-Llama3 [laurençon2024building]. For early-fusion MLLMs, we evaluate Qwen2.5-VL-7B-Instruct [3] and LLaMA-3.2-11B-Vision-Instruct [28]. For commercial MLLMs, we consider Claude-Sonnet-4 [2], Gemini-2.5-Pro [12], and GPT-5 [30].
@@
Datasets and Evaluation Metrics. We evaluate our approach on three benchmarks: MaliciousInstruct [15], AdvBench [64], and HADES [19], containing 100, 520, and 750 malicious instructions, respectively. For textual inputs, we adopt plain malicious prompts without modification. For AdvBench and MaliciousInstruct, the visual input is initialised with either a blank image of RGB $(128,128,128)$ or a panda image [32]. For HADES, we adopt the provided image–instruction pairs (step 5) as initialisation while removing keyword typography to ensure the model focuses on the image content. Regarding commercial models, we test the top 100 instructions from MaliciousInstruct and AdvBench, and the top 20 instructions in HADES spanning five categories. To avoid false positives, we evaluate the attack success rate (ASR) by combining substring matching with LLM-based judgment. Substring matching verifies whether the model refuses to answer the malicious instruction [64], while HarmBenchLLaMA-2-13B-cls [27] determines whether the response is actually harmful. The results of visual jailbreaking attacks against defence techniques are presented in Appendix H.
@@
Setup for FORCE. We set the noise neighborhood to $\eta=4/255$, the regularization strength to $\lambda=0.75$, the scaling factor to $\beta=0.95$, the number of reference samples to $N=10$, and the number of frequency bands to $M=10$. All other settings remain consistent with the baseline PGD to ensure a fair comparison
@@
4.2 Performance Evaluation
To comprehensively evaluate our attack, we examine its cross-model transferability on two different MLLM architectures and API-based MLLMs. From Table 1, we can observe that visual jailbreaking attacks generated by standard PGD exhibit considerable transferability to adapter-based MLLMs, with an average ASR of about 50% and requiring 50 queries per successful attack. For this scenario, our proposed FORCE demonstrates superior performance across all evaluation settings, achieving an average ASR improvement of 12% while reducing the average query cost by over 15%.
Finally, we extend our method to jailbreak commercial MLLMs, which incorporate state-of-the-art alignment techniques and auxiliary safety filters. As shown in Table 1, FORCE can consistently enhance transferability across three mainstream commercial models, achieving an average improvement of 70%. Despite the baseline’s limited capability restricting absolute ASR increases, our method delivers substantial relative improvements and represents a firm step toward practical optimisation-based visual attacks. The case analysis of the FORCE attack can be found in Appendix I.
@@ 4.3 Ablation Study
Zero-shot Transferability. We further evaluate the most stringent zero-shot transferability setting, where only a single query is permitted to jailbreak the target MLLMs. As shown in Table 2 (right), this restrictive scenario leads to a sharp decline in the effectiveness of PGD, which can be attributed to its narrow feasible regions that are difficult to precisely align with the vulnerabilities of target models within a single shot. While this setting also poses challenges for FORCE, its ability to identify a flatter loss landscape increases the likelihood of exploiting target vulnerabilities in a single attempt and improves transferability.
@@
$\checkmark$ - 55.00 (3.8%) 48.46 (4.7%)
- $\checkmark$ 59.00 (11.3%) 44.03 (15.2%)
$\checkmark$ $\checkmark$ 64.00 (20.6%) 39.59 (28.1%)
Optimisation Objectives. To validate the effectiveness of our proposed method in reducing model-specific reliance, we visualise the layers’ feasible regions and the influence of frequency bands. These visualisations follow the same approach described in Section 3.2 and Section 3.3. As presented in Figure 5, it is clear that our method encourages visual jailbreaking attacks to explore broader representations in the early layers, resulting in a smoother loss increase during feature interpolation compared to the baseline in Figure 3. This also drives the attack toward a flatter loss landscape in the input space, thereby improving its resilience to parameter shifts during transfer. We also examine the capability of our method in the spectral domain by analysing the influence of frequency components on attack performance. As depicted in Figure 6, our method reliably mitigates the abnormal reliance on semantically poor information, as evidenced by a more moderate loss change when masking high-frequency informations, and exhibits a natural trend similar to that of non-adversarial images. Both outcomes indicate that FORCE effectively mitigates model-specific reliance, promotes exploration of a flatter loss landscape, and enhances transferability.
Impact of Components. We investigate both the individual and joint contributions of the two components in our algorithm, as presented in Table 3. Our results demonstrate that each component effectively mitigates the improper reliance it is designed to address, thereby yielding a notable improvement in transferability. In particular, the layer-feature regularisation improves transferability by 3.8% and query efficiency by 4.7%, while the spectral-rescaling component yields gains of 11.3% and 15.2%, respectively. Moreover, combining the two components produces a synergistic effect that more thoroughly removes improper reliance, ultimately resulting in an overall performance improvement of 20.6%.
4.4 Generation Costs
We comprehensively evaluate the computational cost and memory footprint of our proposed method. As shown in Table 4, our regularisation term relies only on intermediate variables already produced in the standard PGD attack, incurring negligible additional memory cost, and the multi-sampling process can be computed in parallel with minor computational overhead. More importantly, our objective is to generate transferable visual jailbreaking attacks that assess the vulnerability of different MLLMs via a single-shot process, which is substantially more time-efficient than crafting separate attacks for each model.
Table 4: Comparison of the generation cost of visual jailbreaking attacks. The results are obtained on a single AMD MI250X GPU and averaged over 30 optimisation iterations.
Method Optimisation Time (S) Memory (GB)
PGD 2.17 32.64
FORCE 2.73 36.48
@@ 5 Conclusion
In this work, we investigated the limited transferability of optimisation-based visual jailbreaking attacks and attributed this issue to their reliance on model-specific features in early layers and high-frequency information. This reliance drives the attacks into high-sharpness regions, leaving them vulnerable to parameter shifts during transfer. To address this, we introduced a Feature Over-Reliance CorrEction (FORCE) method, which encourages attacks to explore broader regions in the layer space while rescaling frequency components according to their semantic relevance. By correcting both layer space and spectral domain dependencies, FORCE enables the discovery of flattened feasible regions that enhance cross-model transferability. Extensive experiments demonstrate that our approach provides an important step toward a practical visual red-teaming evaluation. We discuss limitations of our work and future directions in Appendix J.
Acknowledgments
The authors express gratitude to Xiuchuan Li for his helpful feedback. The authors also thank the reviewers and the area chair for their valuable comments. This work was supported by resources provided by the Pawsey Supercomputing Research Centre’s Setonix Supercomputer (https://doi.org/10.48569/18sb-8s43), with funding from the Australian Government and the Government of Western Australia. This work is partly supported by the OpenAI Researcher Access Program. Adel Bibi, Alasdair Paren, and Philip Torr acknowledge the 2025 UK AISI Systemic Safety Grant. Adel Bibi and Philip Torr also acknowledge the UKRI Turing AI Fellowship (EP/W002981/1). Adel Bibi is affiliated with the Institute for Decentralized AI, which is supported by an AI Safety Fund grant. Tongliang Liu is partially supported by the following Australian Research Council projects: FT220100318, DP260102466, DP220102121, LP220100527, LP220200949.
Supplementary Material
@@ Appendix A Feature Interpolation Between Visual Jailbreaking Attacks
We also interpolate features between two different visual jailbreaking attacks generated on LLaVA-v1.5-7B [22], as shown in Figure 7. Consistent with our observation in Section 3.2, we find that feasible regions in later layers are flatter, whereas they become progressively narrower toward earlier layers. Moreover, our results show that in later layers, different jailbreaking examples occupy a shared continuous region, as interpolated attacks consistently succeed in manipulating the source MLLM. In earlier layers, the feasible regions of different attacks become disjoint, as the interpolated features cause them to lose effectiveness. Togetherly, these results reveal that visual attacks tend to rely on model-specific features in earlier layers, leading to small and disjoint feasible regions that fail to generalise across models.
@@
Appendix B Universality of Early-layer Dependency
We further verify our observation about the early-layer dependency on InstructBlip-Vicuna-7B [9]. As shown in Figure 8, visual jailbreaking attacks exhibit progressively narrower feasible regions in shallower layers, indicating that this phenomenon is shared across different model architectures.
Appendix C Universality of High-frequency Dependency
We also verify our observation about the high-frequency dependency on InstructBlip-Vicuna-7B [9]. As shown in Figure 9, the visual attack’s effectiveness also becomes increasingly dependent on high-frequency components, suggesting that this behaviour is independent of the model architecture.
Moreover, we further elucidate this phenomenon by optimising perturbations using only the top 50% of the low- or high-frequency components on LLaVA-v1.5-7B [22]. As shown in Table 5, using high-frequency features provides a more efficient yet superficial shortcut for minimising the loss. Consequently, the tendency to converge toward high-frequency features rather than semantically meaningful content is intrinsic to the optimisation process. Building on this, our spectral feature regularisation mitigates this improper reliance on such shortcuts and improves transferability.
Table 5: Loss curve over the optimisation process.
Iter. PGD low-frequency high-frequency
100 0.492 0.564 1.032
300 0.075 0.532 0.455
500 0.014 0.480 0.267
700 0.013 0.440 0.097
900 0.007 0.363 0.081
Appendix D FORCE Algorithm
@@
3:   Generate spectral-rescaled perturbation via Algorithm 3 $\delta\leftarrow\delta_{rescaled}$
@@
0:  L-layer Network $f_{\mathbf{\theta}}$, input text $\mathbf{x}_{\text{txt}}$, input image $\mathbf{x}_{\text{img}}$, target output $\mathbf{y}$, jailbreaking perturbation $\delta$, number of reference samples NN, noise neighbourhood $\eta$, regularisation strength $\lambda$.
@@
1:  $\lambda_{l}=\lambda\cdot\max\left(1-\left(\tfrac{2\cdot l}{L}\right)^{2},0\right),\qquad l=1,\ldots,L$
@@
4:   Extract layer feature $\mathbf{h}_{\eta_{n},l}=\!\left(f_{\theta,l}(\mathbf{x}_{\text{img}}+\delta+\eta_{n},\mathbf{x}_{\text{txt}})\right),\qquad$ for $l=1,\ldots,L$
@@
7:  Extract layer feature $\mathbf{h}_{jail,l}=\!\left(f_{\theta,l}(\mathbf{x}_{\text{img}}+\delta,\mathbf{x}_{\text{txt}})\right),\qquad$ for $l=1,\ldots,L$
8:  $\ell_{\text{reg}}=\frac{1}{N}\sum_{n=1}^{N}\sum_{l=1}^{L}\left(\lambda_{l}\cdot\frac{\ell_{\text{n}}}{\left\|\mathbf{h}_{\text{jail},l}-\mathbf{h}_{n,l}\right\|_{2}^{2}}\right)$
@@
0:  L-layer Network $f_{\mathbf{\theta}}$, input text $\mathbf{x}_{\text{txt}}$, input image $\mathbf{x}_{\text{img}}$, target output $\mathbf{y}$, jailbreaking perturbation $\delta$, number of frequency bands $M$, scaled factor $\beta$.
@@
1:  $(A,\Phi)\leftarrow\mathrm{FFT}(\delta)$
2:  $\mathcal{B}=\{B_{0},\ldots,B_{M-1}\}\text{ is a partition of }\mathrm{supp}(A),$$\mu(B_{m})=\tfrac{1}{M}\mu(\mathrm{supp}(A))\;\;\forall m,$
@@
8:  $w_{m}=\min\!\left(\beta,\ \frac{\ell_{m-1}}{\ell_{m}}\cdot\beta\right),\qquad m=1,\ldots,M$
9:  $S=\sum_{m=1}^{M}(w_{m}\cdot\mathbbm{1}{B_{m}})$
10:  $A_{\text{rescaled}}=A\odot S$
11:  $\delta_{\text{rescaled}}\leftarrow\mathrm{IFFT}\big(A_{\text{rescaled}}\odot e^{\mathrm{i}\Phi}\big)$
Appendix E Generating Attacks on Different Models
To assess the generality of our approach, we use InstructBLIP-Vicuna-7B [9] as the source MLLM for generating visual jailbreaking attacks. In this setting, all hyperparameters are kept unchanged, except that we set the regularisation strength to $\lambda=0.01$ to adapt to the feature-space scale of InstructBLIP-Vicuna-7B. As shown in Table 6, our method consistently enhances transferability compared with the baseline. For instance, it achieves a 32% improvement in ASR and a 37.2% gain in query efficiency on Idefics3-8B-Llama3 [laurençon2024building]. More importantly, these findings highlight that feature over-reliance is a pervasive issue in optimisation-based visual jailbreaking attacks and demonstrate the general effectiveness of our method.
Table 6: Attack results generated by InstructBLIP-Vicuna-7B on MaliciousInstruct.
Target Model Method ASR ($\uparrow$) Query ($\downarrow$)
Llava-v1.6-mistral-7b [22] PGD 43.00 59.80
FORCE 50.00 52.93